OpenStudy (anonymous):

The circumference (C) of a swimming pool is 47 feet. Which formula can you use to find the radius (r) if you know that C = 2πr ?

OpenStudy (kamibug):

You just need to rearrange that formula so that you can solve for r. Divide both sides by 2pi to isolate r. r=C/2pi

OpenStudy (kamibug):

\(\Huge\bf\color{seagreen}{Welcome~to~Openstudy!}\)

OpenStudy (texaschic101):

C = 2 (pi) (r) 47 = 2 (pi)(r) 47/ ((2)(pi)) = r